FeCo alloy thin film is a fundmental material for the STT-MRAM. It is a FCC structure when disordered and L10 structure when 100% ordered. For the perpendicular magnetic anisotropy, L10 structure with FeCo(100) perpendicular to the surface is preferred. I prepare the thin film by magnetron sputtering. Suppose I should prepare the FeCo alloy with atomic ratio of 1:1 in an ideal case. However, due to poor preparation of co-sputtering, the atom is 5:3 (iron rich FeCo thin film), the structure of Fe rich thin film is noy known. Anonalous X-ray scattering was perform to solve this issue. Three kind of models is proposed. In the first model, the addition Fe is segregated form the order FeCo (1:1) L10 structure, and fill in the grain boundaries. The second model is totally disorder FCC structure (order parameter is zero) and addition Fe is dissolved into a solution. The third model is Fe atom replaces 1/4 part of Co atom positions in the L10 structure.


(a) Please calculate the diffraction intensity ratio of FeCo(001)/FeCo(002) for these three models using 12 KeV X-rays. To simplify the question, we assume the atiomic structure factor is constant (usually, atomic structure factor is a function of q) f_F_e=26, f_C_o=27. In addition, all the FeCo alloys are in cubic structure and the lattice constants are all the same from 0-100% ordering

(b) In order to gain the sensitivity of determining the order parameter, I charge the energy of X-ray form 12 KeV to the K-edge of Fe (7.112 KeV). The atomic structure factor of Fe is changed by adding the dispersive factor of f', which is around 5. Please also calculate the X-ray diffraction intensity ratios of FeCo(001)/FeCo(002) of the FeCo alloy for these three models, Hint: f=f_0-f' and this method called anomalous X-ray scattering method.

(Please, answer All two question please , the type format is the best)
Answer #1

1. For a diffraction intensity the intensity is directly proportional to the structure factor ( F )

I alpha F2

so calculating different value of structure factors for all three models.

1 . for 1st model

F = f( FeCo) ei [ 2pi ( hx + ky + lz ]

also f( FeCo) = fFe + fCo = 26+27 = 53

so F for FeCo ( 001 ) is

F = 53 ei [ 2pi ( h(0) + k(0) + l(1)] = 53 ei [ 2pi( l )] ( h+k+l = odd ) so BCC

for FeCo ( 002 )

F = 53 ei [ 2pi ( h(0) + k( 0) +l (2)] = F = 53 ei 2pi(2l)

2. for 2nd model ( FCC )

the structure value is given as F = f ( 1+eipi ( h+k) + eipi( k+ l )+ eipi( l + h )

for FeCo ( 001) F = 53 ( 1+ eipi(h + k)(0)+ eipi(k+l)(0)+eipi(l + h)(1)

similarly find out for FeCo ( 002) coordinate

3. for third model the Fe atom is replaced by 1/4 th of Co atoms

so fFe = 1/4 x 27 =6.75 so fFeCo = 6.75 + 27 = 33.75

Now the structure value of F can be calculated as the 1st model just by putting the value of fFeCo = 33.75

2. as the value of f' = 5 so,

Fe Co = 53 こ 58. RoHos of fe Co(oo Fe Co (00 2 Re Co Coo Pe Co (002 533-78s 2 84+930.33 > 0,018ilo) fe Co(002) o (oo) Fe Col 002) 2 3 t e 26-14-0.04C 556,